Write the slope-intercept form equation of the line with slope − 1 that passes through the point (0,-5)
Answer by ewatrrr(24785)

 
Hi,
line with slope m = -1 that passes through the point (0,-5)
(y-(-5)) = -1(x-0)
y = -x -5